# Open quantum battery in the background of a three-dimensional rotating black hole

### v1

@@ Abstract
We investigate the charging performance of a quantum battery (QB) coupled to a scalar field under the background of a three-dimensional rotating black hole. It is shown that under Dirichlet boundary conditions, when the QB’s energy level spacing is smaller than the charging amplitude, the black hole rotation enhances the charging performance at finite times, whereas in other parameter regimes, it degrades the charging performance. Notably, as the black hole approaches extremal rotation, charging performance undergoes significant amplification or suppression, depending on the parameter regime. This indicates that the performance of QBs can probe critical properties of black holes. Additionally, regarding the energy flow in QB, it is further demonstrated that the energy extraction from vacuum fluctuations via dissipation, and rotation suppresses the QB’s capacity to extract this energy. Our findings not only advance the relativistic dissipation dynamics of QBs but also propose a novel method to detect black hole rotation and extremal-state transitions.
@@ I Introduction
In recent years, the rapid advancement of thermodynamics and quantum information science has intensified research focus on the quantum thermodynamic properties at their intersection Esposito et al. (2009); Campisi et al. (2011); Landi and Paternostro (2021); Pekola and Karimi (2021). An important example is quantum battery Allahverdyan et al. (2004); Alicki and Fannes (2013), a miniaturized device that leverages quantum effects and quantum operations to achieve advantages in energy storage capacity and charging power Yang et al. (2023); Zhang et al. (2024). To investigate how the charging performance of quantum battery can be enhanced by quantum effects, various quantum battery schemes have been studied from both theoretical and experimental perspectives Zhu et al. (2023); Niu et al. (2024); Hu et al. (2022); Quach et al. (2022); Joshi and Mahesh (2022), such as the Sachdev-Ye-Kitaev quantum battery Rossini et al. (2020); Rosa et al. (2020), the cavity quantum battery Crescente et al. (2020); Lu et al. (2021); Zhang et al. (2019); Dou et al. (2022a); Yang et al. (2024); Wang et al. (2024), and the spin-chain quantum battery Kamin et al. (2020); Grazi et al. (2024); Dou et al. (2022b); Ghosh et al. (2020); Rossini et al. (2019). Most early studies focused on closed quantum battery systems, which consist of a quantum battery and charger isolated from environmental interactions Ferraro et al. (2018); Crescente et al. (2020); Zhang et al. (2019). However, real quantum systems always interact with the environment Breuer and Petruccione (2002), causing relaxation and decoherence. For quantum batteries, these effects can significantly impact the charging performance. Therefore, investigating dissipative charging processes in open quantum systems is of critical importance Barra (2019); Bai and An (2020); Zhao et al. (2021); Gherardini et al. (2020); Mitchison et al. (2021); Xu et al. (2021); Tian et al. (2024); Bhattacharyya et al. (2024); Malavazi et al. (2024); Hu et al. (2025); Tian et al. (2025).
On the other hand, as a result of quantum field scattering by spacetime, the geometric properties of spacetime are typically encoded into the field correlation function, which in turn may influence the dynamical evolution of the detector through its interaction with external quantum field Manzano (2020); Benatti et al. (2003); Benatti and Floreanini (2004); Soares et al. (2022); Yu and Zhang (2008); Yu (2011); Kaplanek and Burgess (2020); Kaplanek and Tjoa (2023). Consequently, the study of open quantum systems within the relativistic framework has received extensive attention, including investigations into quantum field and spacetime properties Tian and Jing (2013); Liu et al. (2024c); Du and Mann (2021); Tian et al. (2015); Tian and Jing (2023), harvesting the vacuum entanglement of quantum field Liu et al. (2025b, 2021); Wu et al. (2024), and so on. In addition, this detector-field interaction model has also been extensively employed to investigate the interplay between the thermodynamics of quantum systems and the superposition property of spacetime Henderson et al. (2020); Martín-Martínez and Perche (2023); Foo et al. (2022).
@@
This work investigates the dissipation dynamics of an open quantum battery under the background of a three-dimensional BTZ black hole Banados et al. (1992). The quantum battery is modeled as a two-level system coupled to the external massless scalar field. By focusing on dissipative dynamics, the charging performance of the battery is analyzed to investigate the interplay between the thermodynamics of quantum systems and gravitational effects. Specifically, we investigate the influence of black hole rotation on the charging performance by varying energy level spacing and charging amplitude. Our results show that when the energy level spacing is smaller than the charging amplitude, dissipation under Dirichlet boundary conditions in the extremal black hole scenario significantly enhances both energy storage capacity and charging power. Furthermore, an analysis of energy flow demonstrates that increased black hole rotation speeds reduce the proportion of energy extracted from vacuum fluctuations.
@@ II Dissipation dynamics of open quantum battery
where $H_{\mathrm{B}}=\frac{\Delta}{2}\sigma_{z}$ and $H_{\mathrm{C}}=-\frac{A}{2}\sigma_{x}$ denote the free Hamiltonian of the quantum batttery and the charger, respectively. $\Delta$ is the energy level spacing between the ground state $\left|g\right\rangle$ and the excited state $\left|e\right\rangle$ of the quantum battery, $\sigma_{m}(m=x,y,z)$ represents the $m\mathrm{-th}$ Pauli matrix, and $A$ denotes an external classical field coupled with the $\sigma_{x}$ component of the quantum battery Binder et al. (2015); Campaioli et al. (2017); Ferraro et al. (2018). Meanwhile, $H_{\mathrm{FI}}=H_{\mathrm{F}}+H_{\mathrm{I}}$ contains two terms: $H_{\mathrm{F}}$ denotes the free Hamiltonian of the scalar field, and $H_{\mathrm{I}}$ denotes the interaction between the quantum battery and the field. Specifically,
$H_{\mathrm{F}}$ $=$ $\int d^{3}\mathbf{k}\omega_{\mathbf{k}}c_{\mathbf{k}}^{\dagger}c_{\mathbf{k}}\frac{dt}{d\tau},$
$H_{\mathrm{I}}$ $=$ ${\lambda}(\sin\theta\sigma_{z}+\cos\theta\sigma_{x})\otimes\phi[x(\tau)],$ (2)
where $c_{\mathbf{k}}^{\dagger}$($c_{\mathbf{k}}$) is the creation (annihilation) operator of the field mode $\mathbf{k}$ with frequency $\omega_{\mathbf{k}}$, $\tau$ denotes the proper time of the quantum battery, $\lambda$ represents the coupling strength between the quantum battery and the field, and $\theta$ characterizes the weight of the quantum battery-field coupling that along the longitudinal $(z)$ and transverse $(x)$ directions. Note that this interaction can capture both decoherence $(\theta=0)$ and pure dephasing $(\theta=\pi/2)$ process. It is the interaction with the field that will lead to the dissipation of the quantum battery, as shown below.
We assume that the initial state of the quantum battery and field is prepared at $\rho_{\mathrm{tot}}(0)=\rho_{\mathrm{B}}(0)\otimes|0\rangle\langle 0|$, where $\rho_{\mathrm{B}}(0)=\frac{1}{2}(\mathrm{I}+\vec{\mathrm{\textbf{b}}}(0)\cdot\vec{\sigma})$, and $\left|0\right\rangle$ represents the vacuum state of the field. At $\tau>0$, we switch on the charger, then the quantum battery begins to be charged. To learn the dynamics of the quantum battery conveniently, we employ a rotation operation (23) on the spin space of the quantum battery. Correspondingly, the rotated time-dependent reduced density matrix of quantum battery reads
@@
At $\tau=0$, the rotated initial reduced density matrix of the quantum battery is given by $\tilde{\rho}_{\mathrm{B}}(0)=\frac{1}{2}(\mathrm{I}+\vec{{\textbf{a}}}(0)\cdot\vec{\sigma})$, with $a_{1}(0)=b_{1}(0)\cos\Theta+b_{3}(0)\sin\Theta$, $a_{2}(0)=b_{2}(0)$, and $a_{3}(0)=b_{3}(0)\cos\Theta-b_{1}(0)\sin\Theta$, and $\Theta=\mathrm{Arccos}\big{(}{\Delta}/{\sqrt{\Delta^{2}+\mathrm{A}^{2}}}\big{)}.$
Under the assumption of weak-coupling limit between the quantum battery and the field, we apply the Born approximation and the Markov approximation Breuer and Petruccione (2002) to derive the equation of motion of the quantum battery. After that, we obtain the reduced density matrix $\tilde{\rho}_{\mathrm{B}}(\tau)$ yields the Lindblad master equation Manzano (2020); Breuer and Petruccione (2002)
@@
[…]
@@
×$\times e^{-\frac{1}{2}(\gamma_{+}+\gamma_{-}+4\gamma_{z})\tau},$
@@
×$\times e^{-\frac{1}{2}(\gamma_{+}+\gamma_{-}+4\gamma_{z})\tau},$
@@ III Quantum battery in rotating BTZ black hole
where $h(r)=-M+\frac{r^{2}}{\ell^{2}}+\frac{J^{2}}{4r^{2}},$ $N_{\varphi}(r)=-\frac{J}{2r^{2}}$, with $M$ being the mass of the black hole, $J$ being its angular momentum, and $\ell$ being the AdS length. Note that the metric (7) is a vacuum solution of Einstein field equations with a negative cosmological constant. The inner and outer horizons (where $h(r)=0$) of the black hole read
@@
Notably, $|J|\leq M\ell$, and in the below analysis, we consider $J\in(0,1)$. When $J/M\ell$ = 0, it indicates a static BTZ black hole. An increase in $J/M\ell$ corresponds to an increase in the rotational velocity of the black hole. The condition $r_{+}=r_{-}$ corresponds to the extremal condition.
To derive the charging dynamics of the quantum battery in the rotating BTZ spacetime, one first needs to known the Wightman function of the scalar field. In this paper, we assume that the scalar field is initially in the Hartle-Hawking vacuum. Since the BTZ spacetime is locally equivalent to the $\mathrm{AdS}_{3}$ spacetime, the Wightman function for the conformal massless scalar field in the rotating BTZ can be represent by Carlip (1995); Lifschytz and Ortiz (1994),
@@
×$\times\bigg{[}\frac{1}{\sqrt{\sigma_{\epsilon}(x,\Gamma^{n}x^{\prime})}}-\frac{\zeta}{\sqrt{\sigma_{\epsilon}(x,\Gamma^{n}x^{\prime})+2}}\bigg{]},$
@@
Substituting the Wightman function (9), together with the quantum battery’s world line (III), into Eq. (II), one can derive
@@
where $T$ is the local KMS temperature, which is the temperature experienced by the quantum battery, satisfying $T=T_{H}/\gamma$, with $\gamma=\frac{\sqrt{R^{2}-r_{+}^{2}}\sqrt{r_{+}^{2}-r_{-}^{2}}}{lr_{+}}$ and $T_{H}=\frac{1}{2\pi\ell}\left(\frac{r_{+}^{2}-r_{-}^{2}}{r_{+}}\right)$. Besides, $P_{v}\left(x\right)$ is the associated Legendre function of the first kind, satisfying $P_{-\frac{1}{2}+i\lambda}=P_{-\frac{1}{2}-i\lambda}$, and the definitions
$\cosh(\alpha_{n}^{\pm})={\pm}4\ell^{2}\pi^{2}T^{2}+(1+4\ell^{2}\pi^{2}T^{2})\cosh\left(\frac{2\pi nr_{+}}{\ell}\right),$ (13)
@@ III.1 Dynamics of the average energy
In this subsection, we investigate how the rotation property of the rotating BTZ black hole affects the average energy stored in the quantum battery during the charging protocol, through dissipation determined by $\gamma_{\pm}$ and $\gamma_{z}$ shown in Eq. (III). In doing so, we assume the quantum battery is initially prepared in the ground state $|g\rangle$, i.e., $b_{1}(0)=0,b_{2}(0)=0,b_{3}(0)=-1$, and consider the decoherence coupling case ($\theta=0$ shown in Eq. (II)). We can obtain the average energy of the quantum battery and the charger Yang et al. (2023)
@@
$\mathcal{K}^{\prime}$ $=$ $\frac{\lambda^{2}\Delta^{2}}{2\Omega^{2}}\sum_{n=-\infty}^{\infty}e^{\pm\frac{i\Omega nr_{-}}{\ell T}}\bigg{[}P_{\frac{i\Omega}{2\pi T}-\frac{1}{2}}\big{(}\cosh\alpha_{n}^{-}\big{)}-\zeta P_{\frac{i\Omega}{2\pi T}-\frac{1}{2}}\big{(}\cosh\alpha_{n}^{+}\big{)}\bigg{]},$
@@
Note that dissipation effects are reflected in the incoherent relaxation rate $\mathcal{K}^{\prime}$ and the dephasing rate $\mathcal{K}$.
The absence of dissipation between the quantum battery and the field is also a scenario we are concerned about, which corresponds to the coupling constant $\lambda$ = 0 in Eq. (II), when the closed quantum battery is driven by a static protocol and all its all energy is supplied by the charger. In this case, the energy stored in the quantum battery is
@@
In addition, there is another special scenario in which the quantum battery interacts with the field in the infinite time limit, i.e., $\tau\to\infty$. In this limit, the average energy stored in the quantum battery (14) is expressed as
@@
In this limit, the average energy of the quantum battery remains unaffected by exponential decay, as indicated by (III.1). This suggests the boundary conditions have no discernible effect on the energy at this stage. Instead, the average energy is solely determined by the angular momentum $J/M\ell$ and relative weighting between the quantum battery’s energy level spacing $\Delta\ell$ and the charger’s amplitude $A\ell$. The effects of these two factors on the average energy stored in a quantum battery in the asymptotic long-time limit are elucidated Fig. 1.
@@
It is demonstrated that the maximum energy stored in a closed quantum battery is solely dependent on the energy level spacing $\Delta\ell$ and charging amplitude $A\ell$. It is noteworthy that no energy can be stored in the closed quantum battery when the ratio of $\Delta\ell$/$A\ell$ = 0. This indicates that although there is a charger providing energy, the energy level spacing $\Delta\ell$ determines the storage capacity, and no energy can be stored in the quantum battery when $\Delta\ell$ = 0. As the $\Delta\ell$/$A\ell$ increase, the maximum energy stored in the closed quantum battery first increases sharply, reaches a peak when the ratio is near one, and then decreases slowly. This means that the charging amplitude $A\ell$ is fixed, the closed quantum battery stores the maximum energy when the energy level spacing $\Delta\ell$ is equal to $A\ell$.
In addition, it is clearly observed that when the energy level spacing $\Delta\ell$ of the quantum battery is larger than the charging amplitude $A\ell$, and as the angular momentum varies, the energy stored in the quantum battery in the long-time limit may exceed the maximum energy of the closed quantum battery. The energy stored in the quantum battery originates from both the charger and dissipation due to rotating BTZ vacuum fluctuations. In other words, the quantum battery can extract energy from the vacuum fluctuations. However, as angular momentum increases beyond approximately 0.9, the stored energy falls below $\langle E_{\mathrm{B}}(\tau)\rangle_{\mathrm{static}}$. This implies that, in the infinite time limit, the black hole’s rotation inhibits the quantum battery’s ability to extract energy from vacuum fluctuations, which in turn weakens its charging performance. When $\Delta\ell$ $<$ $A\ell$, the dissipative quantum battery stores less energy than the closed quantum battery, regardless of angular momentum. This suggests that, in addition to the rotation of the black hole, the weights between $\Delta\ell$ and $A\ell$ affect charging dynamics.

### v3

@@ Abstract
We investigate the charging performance of a quantum battery coupled to a scalar field in the background of a three-dimensional rotating black hole. We show that for Dirichlet boundary conditions, the black hole rotation enhances the charging performance at finite times when the quantum battery’s energy level spacing is smaller than the charging amplitude, whereas it degrades the charging performance in other parameter regimes. Notably, as the black hole approaches extremal rotation, charging performance undergoes significant amplification or suppression, depending on the parameter regime. This indicates that the performance of quantum battery can probe critical properties of black holes. Additionally, regarding the energy flow in quantum battery, it is further demonstrated that the energy extraction from vacuum fluctuations via dissipation, and rotation suppresses the quantum battery’s capacity to extract this energy. Our findings not only advance the relativistic dissipation dynamics of quantum battery but also propose a novel method to detect black hole rotation and extremal-state transitions.
@@ I Introduction
In recent years, the rapid advancement of thermodynamics and quantum information science has intensified research focus on the quantum thermodynamic properties at their intersection Esposito et al. (2009); Campisi et al. (2011); Landi and Paternostro (2021); Pekola and Karimi (2021). An important example is quantum battery Allahverdyan et al. (2004); Alicki and Fannes (2013); Yang et al. (2023), a miniaturized device that leverages quantum effects and quantum operations to achieve advantages in energy storage capacity and charging power. To investigate how the charging performance of quantum battery can be enhanced by quantum effects, various quantum battery schemes have been studied from both theoretical and experimental perspectives Zhu et al. (2023); Niu et al. (2024); Quach et al. (2022); Joshi and Mahesh (2022), such as the Sachdev-Ye-Kitaev quantum battery Rossini et al. (2020); Rosa et al. (2020), the cavity quantum battery Crescente et al. (2020); Wang et al. (2024), and the spin-chain quantum battery Grazi et al. (2024); Rossini et al. (2019); Liu et al. (2025a). Most early studies focused on closed quantum battery systems, which consist of a quantum battery and charger isolated from environmental interactions Crescente et al. (2020); Zhang et al. (2019). However, real quantum systems always interact with the environment Breuer and Petruccione (2002), causing relaxation and decoherence. For quantum battery, these effects can significantly impact the charging performance. Therefore, investigating dissipative charging processes in open quantum systems is of critical importance Barra (2019); Zhao et al. (2021); Bhattacharyya et al. (2024); Malavazi et al. (2024); Hu et al. (2025); Ghosh et al. (2025); Ahuja et al. (2025).
On the other hand, as a result of quantum field scattering by spacetime, the geometric and quantum properties of spacetime are typically encoded into the field correlation function, which in turn may influence the dynamical evolution of the detector through its interaction with the external quantum field Benatti et al. (2003); Benatti and Floreanini (2004); Soares et al. (2022); Yu (2011); Kaplanek and Tjoa (2023); Wu et al. (2025a); Moustos and Abah (2025a, b). Consequently, the interplay between quantum fields and spacetime geometry has attracted considerable attention, ranging from open quantum system approaches in relativistic settings to broader investigations in quantum gravity and quantum information theory, including studies on field correlations and spacetime structure Tian and Jing (2013); Liu et al. (2024a); Du and Mann (2021); Tian et al. (2015); Tian and Jing (2023); Wu et al. (2023, 2024, 2025b); Li et al. ; Liu et al. (2024b, 2025b, 2025c), as well as vacuum entanglement harvesting and related phenomena Liu et al. (2025d); Bueley et al. (2022); Liu et al. (2025e); Yang et al. (2025); Lopez-Raven et al. (2025); Wang et al. (2025); Ng et al. (2022), among others. In addition, this detector-field interaction model has also been extensively employed to investigate the interplay between the thermodynamics of quantum systems and the superposition property of spacetime Henderson et al. (2020); Tang et al. (2025); Foo et al. (2022, 2023). In addition, a quantum battery Mukherjee et al. (2024); Liu et al. (2025f) modeled as an Unruh-DeWitt detector was examined to investigate how the vacuum fluctuations of a massless scalar field affected the battery’s charging performance in the background of a static (2+1)-dimensional Bañados-Teitelboim-Zanelli (BTZ) black hole spacetime Tian et al. (2025). However, the current understanding of how frame-dragging induced by a rotating black hole influences the performance of a quantum battery remains limited. A black hole possessing nonzero angular momentum alters the structure of the surrounding quantum vacuum, making it qualitatively distinct from that associated with a non-rotating black hole. Studies have indicated that rotation can significantly amplify both the entanglement harvested by Unruh-DeWitt detectors Robbins et al. (2021) and the strength of the weak anti-Hawking effect Robbins and Mann (2022).
@@
Motivated by the above considerations, we study the dissipative dynamics of an open quantum battery in the background of a three-dimensional rotating BTZ black hole. The quantum battery is modeled as a two-level system coupled to the external massless scalar field. By focusing on dissipative dynamics, the charging performance of the battery is analyzed to investigate the interplay between the thermodynamics of quantum systems and gravitational effects. Specifically, we explore how the rotation of black holes affects the charging dynamics and overall performance of quantum battery compared with the static BTZ case. Our results show that when the energy level spacing is smaller than the charging amplitude, dissipation for Dirichlet boundary conditions in the extremal black hole scenario significantly enhances both energy storage capacity and charging power. Furthermore, an analysis of energy flow demonstrates that increased black hole rotation speeds reduce the proportion of energy extracted from vacuum fluctuations.
@@ II Dissipation dynamics of open quantum battery
where $H_{\mathrm{B}}=\frac{\Delta}{2}\sigma_{z}$ and $H_{\mathrm{C}}=-\frac{A}{2}\sigma_{x}$ denote the free Hamiltonian of the quantum batttery and the charger, respectively. $\Delta$ is the energy level spacing between the ground state $\left|g\right\rangle$ and the excited state $\left|e\right\rangle$ of the quantum battery, $\sigma_{m}(m=x,y,z)$ represents the $m\mathrm{-th}$ Pauli matrix, and $A$ denotes an external classical field coupled with the $\sigma_{x}$ component of the quantum battery Binder et al. (2015); Campaioli et al. (2017); Ferraro et al. (2018). Meanwhile, $H_{\mathrm{FI}}=H_{\mathrm{F}}+H_{\mathrm{I}}$ contains two terms: $H_{\mathrm{F}}=\int d^{3}\mathbf{k}\omega_{\mathbf{k}}c_{\mathbf{k}}^{\dagger}c_{\mathbf{k}}\frac{dt}{d\tau}$ denotes the free Hamiltonian of the scalar field, and $H_{\mathrm{I}}={\lambda}(\sin\theta\sigma_{z}+\cos\theta\sigma_{x})\otimes\phi[x(\tau)]$ denotes the interaction between the quantum battery and the field. Here, $c_{\mathbf{k}}^{\dagger}$($c_{\mathbf{k}}$) is the creation (annihilation) operator of the field mode $\mathbf{k}$ with frequency $\omega_{\mathbf{k}}$, $\tau$ denotes the proper time of the quantum battery, $\lambda$ represents the coupling strength between the quantum battery and the field, and $\theta$ characterizes the weight of the quantum battery-field coupling that along the longitudinal $(z)$ and transverse $(x)$ directions. Note that this interaction can capture both decoherence $(\theta=0)$ and pure dephasing $(\theta=\pi/2)$ process. It is the interaction with the field that will lead to the dissipation of the quantum battery, as shown below.
We assume that the initial state of the quantum battery and field is prepared at $\rho_{\mathrm{tot}}(0)=\rho_{\mathrm{B}}(0)\otimes|0\rangle\langle 0|$, where $\rho_{\mathrm{B}}(0)=\frac{1}{2}(\mathrm{I}+\vec{\mathrm{\textbf{b}}}(0)\cdot\vec{\sigma})$, and $\left|0\right\rangle$ represents the vacuum state of the field. At $\tau=0^{+}$, we switch on the charger, then the quantum battery begins to be charged. To learn the dynamics of the quantum battery conveniently, we employ a rotation operation in Eq.(22) on the spin space of the quantum battery. Correspondingly, the rotated time-dependent reduced density matrix of quantum battery reads
@@
At $\tau=0$, the rotated initial reduced density matrix of the quantum battery is given by $\tilde{\rho}_{\mathrm{B}}(0)=\frac{1}{2}(\mathrm{I}+\vec{{\textbf{a}}}(0)\cdot\vec{\sigma})$, with $a_{1}(0)=b_{1}(0)\cos\Theta+b_{3}(0)\sin\Theta$, $a_{2}(0)=b_{2}(0)$, and $a_{3}(0)=b_{3}(0)\cos\Theta-b_{1}(0)\sin\Theta$, and $\Theta=\mathrm{Arccos}\big({\Delta}/\Omega\big),$ and $\Omega={\sqrt{\Delta^{2}+\mathrm{A}^{2}}}$.
For the assumption of weak-coupling limit between the quantum battery and the field, we apply the Born approximation and the Markov approximation Breuer and Petruccione (2002) to derive the equation of motion of the quantum battery. After that, we obtain the reduced density matrix $\tilde{\rho}_{\mathrm{B}}(\tau)$ yields the Lindblad master equation Manzano (2020); Breuer and Petruccione (2002)
@@
$\gamma_{\pm}$ $=$ $\lambda^{2}\cos^{2}(\theta-\Theta)\int_{-\infty}^{+\infty}d\Delta\tau e^{\mp i\Omega\Delta\tau}W(x(\tau),x(\tau^{\prime})),$
$\gamma_{z}$ $=$ $\lambda^{2}\sin^{2}(\theta-\Theta)\int_{-\infty}^{+\infty}d\Delta\tau W(x(\tau),x(\tau^{\prime})),$ (4)
@@
$\times e^{-\frac{1}{2}(\gamma_{+}+\gamma_{-}+4\gamma_{z})\tau},$
@@
$\times e^{-\frac{1}{2}(\gamma_{+}+\gamma_{-}+4\gamma_{z})\tau},$
@@ III Quantum battery in rotating BTZ black hole
where $h(r)=-M+\frac{r^{2}}{\ell^{2}}+\frac{J^{2}}{4r^{2}},$ $N_{\varphi}(r)=-\frac{J}{2r^{2}}$, with $M$ being the mass of the black hole, $J$ being its angular momentum, and $\ell$ being the AdS length. Note that the metric in Eq.(6) is a vacuum solution of Einstein field equations with a negative cosmological constant. The inner and outer horizons (where $h(r)=0$) of the black hole read
@@
Notably, $|J|\leq M\ell$, and in the below analysis, we consider $J\in[0,1)$. When $J/M\ell$ = 0, it indicates a static BTZ black hole. An increase in $J/M\ell$ corresponds to an increase in the rotational velocity of the black hole. The condition $r_{+}=r_{-}$ corresponds to the extremal condition.
To derive the charging dynamics of the quantum battery in the rotating BTZ spacetime, one first needs to known the Wightman function of the scalar field. In this paper, we assume that the scalar field is initially in the Hartle-Hawking vacuum. Since the BTZ spacetime is locally equivalent to the $\mathrm{AdS}_{3}$ spacetime, the Wightman function for the conformal massless scalar field in the rotating BTZ can be represent by Carlip (1995); Lifschytz and Ortiz (1994)
@@
$\times\bigg[\frac{1}{\sqrt{\sigma_{\epsilon}(x,\Gamma^{n}x^{\prime})}}-\frac{\zeta}{\sqrt{\sigma_{\epsilon}(x,\Gamma^{n}x^{\prime})+2}}\bigg],$
@@
Substituting the Wightman function in Eq. (8), together with the quantum battery’s world line in Eq. (III), into Eq. (II), one can derive
@@
where $T$ is the local KMS temperature, which is the temperature experienced by the quantum battery, satisfying $T=T_{H}/\gamma$, with $\gamma=\frac{\sqrt{R^{2}-r_{+}^{2}}\sqrt{r_{+}^{2}-r_{-}^{2}}}{lr_{+}}$ and $T_{H}=\frac{1}{2\pi\ell^{2}}\left(\frac{r_{+}^{2}-r_{-}^{2}}{r_{+}}\right)$. Besides, $P_{v}\left(x\right)$ is the associated Legendre function of the first kind, satisfying $P_{-\frac{1}{2}+i\lambda}=P_{-\frac{1}{2}-i\lambda}$, and the definitions
$\cosh(\alpha_{n}^{\pm})={\pm}4\ell^{2}\pi^{2}T^{2}+(1+4\ell^{2}\pi^{2}T^{2})\cosh\left(\frac{2\pi nr_{+}}{\ell}\right)$ (12)
@@ III.1 Dynamics of the average energy
In this subsection, we investigate how the rotation property of the rotating BTZ black hole affects the average energy stored in the quantum battery during the charging protocol, through dissipation determined by $\gamma_{\pm}$ and $\gamma_{z}$ shown in Eq. (III). In doing so, we assume the quantum battery is initially prepared in the ground state $|g\rangle$, i.e., $b_{1}(0)=0,b_{2}(0)=0,b_{3}(0)=-1$, and consider the decoherence coupling case ($\theta=0$ shown in $H_{I}$). We can obtain the average energy of the quantum battery and the charger Yang et al. (2023)
@@
$\mathcal{K}^{\prime}$ $=$ $\frac{\lambda^{2}\Delta^{2}}{2\Omega^{2}}\sum_{n=-\infty}^{\infty}e^{\frac{i\Omega nr_{-}}{\ell T}}\bigg[P_{\frac{i\Omega}{2\pi T}-\frac{1}{2}}\big(\cosh\alpha_{n}^{-}\big)-\zeta P_{\frac{i\Omega}{2\pi T}-\frac{1}{2}}\big(\cosh\alpha_{n}^{+}\big)\bigg],$
@@
Note that dissipation effects are reflected in the incoherent relaxation rate $\mathcal{K}^{\prime}$ and the dephasing rate $\mathcal{K}$. Specifically, Eq. (15) involves an infinite series $n$; to ensure the numerical accuracy of the results presented, we performed numerical convergence tests. The results indicate that the numerical values have converged and become stable when the summation is truncated at $n=5$. Including additional terms beyond $n=5$ does not affect the reported results but increases computational cost. Therefore, in the following section, we set the truncation order $n=5$ when calculating the battery’s average energy.
The absence of dissipation between the quantum battery and the field is also a scenario we are concerned about, which corresponds to the coupling constant $\lambda$ = 0 in $H_{I}$, when the closed quantum battery is driven by a static protocol and all its all energy is supplied by the charger. In this case, the energy stored in the quantum battery is
@@
In addition, there is another special scenario in which the quantum battery interacts with the field in the infinite time limit, i.e., $\tau\to\infty$. In this limit, the average energy in Eq. (13) stored in the quantum battery is expressed as
@@
Eq. (17) agrees with the corresponding expression in the static BTZ case Tian et al. (2025). This agreement arises because, in the long-interaction limit, the average energy given in Eq. (13) is no longer affected by the exponential decay term in Eq. (III.1). This also suggests the boundary conditions have no discernible effect on the energy at this stage. Instead, the average energy is solely determined by local KMS temperature $T$ of rotating BTZ black hole and relative weighting between the quantum battery’s energy level spacing $\Delta\ell$ and the charger’s amplitude $A\ell$. Since the redshift factor $\gamma$ and the Hawking temperature $T_{H}$, which together determine the local temperature $T$, indirectly depend on the black hole’s angular momentum $J/Ml$ (see Eq. (7)), we plot the average energy $\langle E_{\mathrm{B}}(\infty)\rangle$ as a function of $J/Ml$ and the ratio $\Delta\ell/A\ell$ in Fig. 1 to investigate the impact of rotation on quantum battery dynamics.
@@
The figure shows that, in the absence of dissipation (i.e., the quantum battery is not coupled to the vacuum field), the maximum energy $\langle E_{\mathrm{B}}(\tau)\rangle_{\mathrm{static}}$ stored in the closed quantum battery depends only on the energy level spacing $\Delta\ell$ and the charging amplitude $A\ell$. And no energy is stored in the battery when the ratio of $\Delta\ell$$A\ell$ = 0. The reason is that, even when the battery is charged, the energy level spacing $\Delta\ell$ determines its storage capacity. When $\Delta\ell$=0, the battery has zero storage capacity. As the $\Delta\ell$$A\ell$ increase, the maximum energy $\langle E_{\mathrm{B}}(\tau)\rangle_{\mathrm{static}}$ stored in the closed quantum battery first increases sharply, reaches a peak when the ratio of $\Delta\ell$$A\ell$ is near one, and then decreases slowly. This implies that the charging amplitude $A\ell$ is fixed, the quantum battery stores its maximum energy when $\Delta\ell=A\ell$, which equals $A/2$.
We find that in the long-time limit quantum battery interacting with the vacuum field may on average store more energy than the closed quantum battery. This behavior occurs when both $\Delta\ell>A\ell$ and the angular momentum $J/M\ell$ << 0.95. At this point, the energy stored in the quantum battery will originate not only from the charger but also from vacuum fluctuations via dissipative processes. Specifically, we find that when the angular momentum $J/M\ell=0$, i.e., for a non-rotating black hole, the quantum battery extracts the maximum possible energy from vacuum fluctuations. However, as the angular momentum $J/Ml$ increases (that is, as the black hole’s rotational speed rises) the energy $\langle E_{\mathrm{B}}(\infty)\rangle$ gradually decreases. This implies that, in the absence of boundary conditions, the nonzero angular momentum modifies the structure of the vacuum field around the black hole, thereby reducing the quantum battery’s ability to extract energy from vacuum fluctuations. Moreover, the black hole’s rotation further amplifies this effect.